# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v3

@@ A.3 Proof of Theorem 2
$L(\theta)-l(\theta)$ $=\big{[}\widehat{B}_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)\big{]}+\big{[}\widehat{B}_{\breve{\gamma}}(\theta)-B_{\breve{\gamma}}(\theta)\big{]}$
$\quad-\bigg{[}\frac{1}{2}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)\bigg{]}-\bigg{[}\frac{1}{2}s_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)s_{\breve{\gamma}}(\theta)-B_{\breve{\gamma}}(\theta)\bigg{]}$
$+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$
Moreover, from Lemma 1,
$\frac{1}{2}s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)-B_{\breve{\alpha}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},\quad\text{and}\quad\frac{1}{2}s_{\breve{\gamma}}^{\prime}\left(\theta\right)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)s_{\breve{\gamma}}(\theta)-B_{\breve{\gamma}}(\theta)=o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}.$
Thus, combining this with (A.16) yields the desired result.
A.4 Proof of Corollary 1
Consider a mean value expansion of $\triangledown_{\theta}L(\widehat{\theta}_{L})$ around $\theta_{0}:$
$0=\triangledown_{\theta}L(\widehat{\theta}_{L})$ $=\triangledown_{\theta}L(\theta_{0})+\triangledown_{\theta\theta^{\prime}}L(\overline{\theta})(\widehat{\theta}_{L}-\theta_{0}),$ (A.21)
@@ A.4 Proof of Corollary 1
$0=$ $\triangledown_{\theta}l(\theta_{0})+\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})(\widehat{\theta}_{L}-\theta_{0})+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)}$ (A.22)
@@
$\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})$ $=-\big{(}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (A.23)
Given the conditional cross-sectional independence and the conditional strong mixing condition (Assumption 1), (A.23) shows that
$\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}\left(0,\Sigma^{-1}\Omega\Sigma^{-1}\right)$ (A.24)
where $\mathbb{E}[\triangledown_{\theta}l(\theta_{0})]=0,$ and
$\Omega=\lim_{N,T\rightarrow\infty}(NT)\mathbb{E}[\triangledown_{\theta}l(\theta_{0})\triangledown_{\theta^{\prime}}l(\theta_{0})].$
Under Assumption 2, the information matrix equality ($\Sigma^{-1}=\Omega$) holds, where the asymptotic variance in (A.24) simplifies to $\Sigma$.
A.5 Proof of Lemma A.3
@@
Under Assumptions 1 and 2, and under the null hypothesis $H_{0}:R(\theta_{0})=0$, as $N$ and $T\to\infty,$
$\widehat{\theta}_{R}-\widehat{\theta}_{L}=\Xi(\theta_{0})\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
@@
$\Xi(\theta_{0}):=\big{(}\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}[J(\theta_{0})]^{\prime}\Big{(}J(\theta_{0})\big{(}\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}[J(\theta_{0})]^{\prime}\Big{)}^{-1}J(\theta_{0})\big{(}\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}.$
@@
$\big{(}\widehat{\theta}_{R},\widehat{\lambda}\big{)}=\arg\max_{\theta\in\Theta,\lambda\in\mathbb{R}}\big{[}L(\theta)+\lambda^{\prime}R(\theta)\big{]}$
@@
Following a similar proof as in Corollary 1, we have $\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}=O_{\mathbb{P}}\left(1\right).$ Furthermore, $\sqrt{NT}\widehat{\lambda}=O_{\mathbb{P}}(1)$ as implied by (A.25) since $\sqrt{NT}\triangledown_{\theta}L(\theta_{0})=O_{\mathbb{P}}(1)$ and $J(\theta_{0})$ is of full row rank. By Theorem 2, we express the mean value expansion of $\triangledown_{\theta}L(\widehat{\theta}_{R})$ around $\theta_{0}$ as
$\triangledown_{\theta}L(\widehat{\theta}_{R})$ $=\triangledown_{\theta}L(\theta_{0})+\triangledown_{\theta\theta^{\prime}}L(\overline{\theta})\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}$
$=\triangledown_{\theta}l(\theta_{0})+\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
@@
$\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}+[J(\theta_{0})]^{\prime}\sqrt{NT}\widehat{\lambda}=-\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (A.27)
Under $H_{0},$ from (A.26), we also obtain
@@
$=\sqrt{NT}R(\theta_{0})+J(\theta_{0})\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}+o_{\mathbb{P}}(1)$
$=J(\theta_{0})\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}+o_{\mathbb{P}}(1).$ (A.28)
@@
$\sqrt{NT}\big{(}\widehat{\theta}_{R}-\theta_{0}\big{)}=-\left(\big{(}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}-\Xi\left(\theta_{0}\right)\right)\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}\left(1\right).$ (A.29)
@@
A.6 Proof of Corollary 2
Let $\widehat{\xi}:=-NT\triangledown_{\theta^{\prime}}l(\theta_{0})\Xi\left(\theta_{0}\right)\triangledown_{\theta}l(\theta_{0}).$ We first show that
@@ A.6 Proof of Corollary 2
${L}(\widehat{\theta}_{R})={L}(\widehat{\theta}_{L})+\triangledown_{\theta^{\prime}}{L}(\widehat{\theta}_{L})\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}+\frac{1}{2}\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}^{\prime}\triangledown_{\theta\theta^{\prime}}L(\theta^{\ast})\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}$
where each component of $\theta^{\ast}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\widehat{\theta}_{L}$. Noting that $\triangledown_{\theta}{L}(\widehat{\theta}_{L})=0,$ we obtain
$\widehat{\xi}_{LR}=-NT\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}^{\prime}\triangledown_{\theta\theta^{\prime}}L(\theta^{\ast})\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}.$
Since $\triangledown_{\theta\theta^{\prime}}L(\theta^{\ast})-\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]=o_{\mathbb{P}}(1),$ by Lemma A.3, we further conclude that
@@
$\triangledown_{\theta}L(\widehat{\theta}_{R})$ $=\triangledown_{\theta}L(\widehat{\theta}_{L})+\triangledown_{\theta\theta^{\prime}}L(\dot{\theta})\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}$
$=\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
where each component of $\dot{\theta}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\widehat{\theta}_{L}.$ Thus,
$\widehat{\xi}_{LM}$ $=-NT\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}^{\prime}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})][\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{R})]^{-1}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}+o_{\mathbb{P}}\left(1\right)$
$=-NT\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}^{\prime}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]\big{(}\widehat{\theta}_{R}-\widehat{\theta}_{L}\big{)}+o_{\mathbb{P}}\left(1\right)$
@@
$R(\widehat{\theta}_{L})=R(\widehat{\theta}_{R})+J(\ddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})=J(\ddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})$
[…]
$\widehat{\xi}_{Wald}$ $=-NT(\widehat{\theta}_{L}-\widehat{\theta}_{R})^{\prime}J(\ddot{\theta})^{\prime}\left(J(\widehat{\theta}_{L})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}J(\widehat{\theta}_{L})^{\prime}\right)^{-1}J(\ddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})$
@@
Next we show
$\widehat{\xi}\overset{\mathbb{D}}{\longrightarrow}\chi^{2}\left(r\right).$ (A.31)
Since $\sqrt{NT}\triangledown_{\theta}l(\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}(0,\Sigma),$ by the delta method, we have
$J(\theta_{0})\big{(}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\big{)}^{-1}\sqrt{NT}\triangledown_{\theta}l(\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}\big{(}0,J(\theta_{0})\Sigma J(\theta_{0})^{\prime}\big{)}.$
This implies (A.31). Therefore, the desired result follows from (A.30) and (A.31). The proof is complete.
B Additional Simulation Results

### v4

@@ A Proofs
$L(\theta)-l(\theta)$ $=\left[\widehat{B}_{\alpha}(\theta)-B_{\alpha}(\theta)\right]+\left[\widehat{B}_{\gamma}(\theta)-B_{\gamma}(\theta)\right]$
$\quad\quad-\left[\frac{1}{2}\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)-B_{\alpha}(\theta)\right]-\left[\frac{1}{2}\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)-B_{\gamma}(\theta)\right]$
$\quad\quad+o_{\mathbb{P}}(1/\sqrt{NT}).$
As from Lemma A.1, we have
$\frac{1}{2}\widetilde{s}_{\alpha}^{\prime}(\theta)H_{\alpha\alpha^{\prime}}^{\ast}(\theta)\widetilde{s}_{\alpha}(\theta)-B_{\alpha}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),\qquad\frac{1}{2}\widetilde{s}_{\gamma}^{\prime}(\theta)H_{\gamma\gamma^{\prime}}^{\ast}(\theta)\widetilde{s}_{\gamma}(\theta)-B_{\gamma}(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}).$
The conclusion follows by combining this and (A.19).
We continue to establishing the asymptotitc properties of $\{\widehat{\theta}_{L},\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}\}$. We first prove Corollary 1 below which states the consistency and asymptotic normality of $\widehat{\theta}_{L}$.
Proof of Corollary 1. Consider a mean value expansion of $\triangledown_{\theta}L(\widehat{\theta}_{L})$ around $\theta_{0}$:
$0=\triangledown_{\theta}L(\widehat{\theta}_{L})=\triangledown_{\theta}L(\theta_{0})+\triangledown_{\theta\theta^{\prime}}L(\overline{\theta})(\widehat{\theta}_{L}-\theta_{0}),$
@@
$0=\triangledown_{\theta}l(\theta_{0})+\triangledown_{\theta\theta^{\prime}}l(\overline{\theta})(\widehat{\theta}_{L}-\theta_{0})+o_{\mathbb{P}}(1/\sqrt{NT})$ (A.24)
@@
$\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})=-\left\{\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\right\}^{-1}\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (A.25)
Given the conditional cross-sectional independence and the conditional strong mixing condition (Assumption 1), (A.25) leads to
$\left[\Sigma^{-1}\Omega\Sigma^{-1}\right]^{-1/2}\sqrt{NT}(\widehat{\theta}_{L}-\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}\left(0,\mathbb{I}_{K}\right)$ (A.26)
as $N,T\rightarrow\infty$, where $\mathbb{E}[\triangledown_{\theta}l(\theta_{0})]=0$ and
$\Sigma:=-\mathbb{E}\left[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})\right],\qquad\Omega:=\left(NT\right)\mathbb{E}[\triangledown_{\theta}l(\theta_{0})\triangledown_{\theta^{\prime}}l(\theta_{0})].$
Since we consider a likelihood model, $\Sigma^{-1}\Omega\Sigma^{-1}$ in (A.26) simplifies to $\Sigma^{-1}$ under the information matrix equality ($\Sigma=\Omega$).
To show the asymptotic distribution of $\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}$, we find the following lemma useful.
@@
Under Assumptions 1 and 2, and under the null hypothesis $H_{0}:R(\theta_{0})=0$, as $N$ and $T\rightarrow\infty,$
$\widehat{\theta}_{R}-\widehat{\theta}_{L}=\Xi(\theta_{0})\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1/\sqrt{NT}),$
@@
$\Xi(\theta_{0}):=\left\{\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\right\}^{-1}[J(\theta_{0})]^{\prime}\left\{J(\theta_{0})\left(\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\right)^{-1}[J(\theta_{0})]^{\prime}\right\}^{-1}J(\theta_{0})\left\{\mathbb{E[}\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\right\}^{-1}.$
@@
$\left\{\widehat{\theta}_{R},\widehat{\lambda}\right\}=\arg\max_{\theta\in\Theta,\lambda\in\mathbb{R}}\left[L(\theta)+\lambda^{\prime}R(\theta)\right]$
@@
Following a similar proof as in Corollary 1, we have $\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})=O_{\mathbb{P}}\left(1\right)$. Furthermore, $\sqrt{NT}\widehat{\lambda}=O_{\mathbb{P}}(1)$ as implied by (A.27) since $\sqrt{NT}\triangledown_{\theta}L(\theta_{0})=O_{\mathbb{P}}(1)$ and $J(\theta_{0})$ is of full row rank. By Theorem 2, a mean value expansion of $\triangledown_{\theta}L(\widehat{\theta}_{R})$ around $\theta_{0}$ is
$\triangledown_{\theta}L(\widehat{\theta}_{R})$ $=\triangledown_{\theta}L(\theta_{0})+\triangledown_{\theta\theta^{\prime}}L(\overline{\theta})(\widehat{\theta}_{R}-\theta_{0})$
$=\triangledown_{\theta}l(\theta_{0})+\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})](\widehat{\theta}_{R}-\theta_{0})+o_{\mathbb{P}}(1/\sqrt{NT}),$
@@
$\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})+[J(\theta_{0})]^{\prime}\sqrt{NT}\widehat{\lambda}=-\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}(1).$ (A.29)
Under $H_{0}$ and from (A.28), we also obtain
@@
$=\sqrt{NT}R(\theta_{0})+J(\theta_{0})\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})+o_{\mathbb{P}}(1)$
$=J(\theta_{0})\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})+o_{\mathbb{P}}(1).$ (A.30)
@@
$\sqrt{NT}(\widehat{\theta}_{R}-\theta_{0})=-\left(\{\mathbb{E}[\triangledown_{\theta\theta^{\prime}}l(\theta_{0})]\}^{-1}-\Xi\left(\theta_{0}\right)\right)\sqrt{NT}\triangledown_{\theta}l(\theta_{0})+o_{\mathbb{P}}\left(1\right).$ (A.31)
@@
Finally, we proof the asymptotic distribution of $\widehat{\xi}_{LR},\widehat{\xi}_{LM},\widehat{\xi}_{Wald}$ below.
Proof of Corollary 2. Let $\widehat{\xi}:=-NT\triangledown_{\theta^{\prime}}l(\theta_{0})\Xi\left(\theta_{0}\right)\triangledown_{\theta}l(\theta_{0})$ for $\Xi\left(\theta_{0}\right)$ defined in Lemma A.4. We first show that
@@
${L}(\widehat{\theta}_{R})={L}(\widehat{\theta}_{L})+\triangledown_{\theta^{\prime}}{L}(\widehat{\theta}_{L})(\widehat{\theta}_{R}-\widehat{\theta}_{L})+\frac{1}{2}(\widehat{\theta}_{R}-\widehat{\theta}_{L})^{\prime}\triangledown_{\theta\theta^{\prime}}L(\dot{\theta})(\widehat{\theta}_{R}-\widehat{\theta}_{L})$
where each component of $\dot{\theta}$ lies between the corresponding components of $\widehat{\theta}_{R}$ and $\widehat{\theta}_{L}$. Noting that $\triangledown_{\theta}{L}(\widehat{\theta}_{L})=0$, we obtain
$\widehat{\xi}_{LR}=-NT(\widehat{\theta}_{R}-\widehat{\theta}_{L})^{\prime}\triangledown_{\theta\theta^{\prime}}L(\dot{\theta})(\widehat{\theta}_{R}-\widehat{\theta}_{L}).$
Since $\triangledown_{\theta\theta^{\prime}}L(\dot{\theta})-\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})]=o_{\mathbb{P}}(1)$, by Lemma A.4, we further conclude that
@@
$\triangledown_{\theta}L(\widehat{\theta}_{R})$ $=\triangledown_{\theta}L(\widehat{\theta}_{L})+\triangledown_{\theta\theta^{\prime}}L(\ddot{\theta})(\widehat{\theta}_{R}-\widehat{\theta}_{L})$
$=\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})](\widehat{\theta}_{R}-\widehat{\theta}_{L})+o_{\mathbb{P}}(1/\sqrt{NT}),$
[…]
$\widehat{\xi}_{LM}$ $=-NT(\widehat{\theta}_{R}-\widehat{\theta}_{L})^{\prime}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})][\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{R})]^{-1}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})](\widehat{\theta}_{R}-\widehat{\theta}_{L})+o_{\mathbb{P}}\left(1\right)$
$=-NT(\widehat{\theta}_{R}-\widehat{\theta}_{L})^{\prime}\mathbb{E}[\triangledown_{\theta\theta^{\prime}}L(\theta_{0})](\widehat{\theta}_{R}-\widehat{\theta}_{L})+o_{\mathbb{P}}\left(1\right)$
@@
$R(\widehat{\theta}_{L})=R(\widehat{\theta}_{R})+J(\dddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})=J(\dddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})$
where each component of $\dddot{\theta}$ lies between the corresponding components of $\widehat{\theta}_{L}$ and $\widehat{\theta}_{R}$. Since both $\widehat{\theta}_{L}$ and $\widehat{\theta}_{R}$(under $H_{0}$) are consistent estimators for $\theta_{0},$ we obtain that
$\widehat{\xi}_{Wald}$ $=-NT(\widehat{\theta}_{L}-\widehat{\theta}_{R})^{\prime}J(\dddot{\theta})^{\prime}\left\{J(\widehat{\theta}_{L})[\triangledown_{\theta\theta^{\prime}}L(\widehat{\theta}_{L})]^{-1}J(\widehat{\theta}_{L})^{\prime}\right\}^{-1}J(\dddot{\theta})(\widehat{\theta}_{L}-\widehat{\theta}_{R})$
@@
Next, since $\Sigma^{-1/2}\sqrt{NT}\triangledown_{\theta}l(\theta_{0})\overset{\mathbb{D}}{\longrightarrow}\mathcal{N}(0,\mathbb{I}_{K})$, applying the delta method delivers $\widehat{\xi}\overset{\mathbb{D}}{\longrightarrow}\chi^{2}\left(r\right)$. The desired results follows from here.
B Extra Simulation Results